A student has 10 seconds left before he is considered late in his 7:oo a.m. class. The classroom is located on the second floor, near the stairway. The 10-second margin starts as he reaches the stairway, which is about 6m high. If the student weighs 500N and he exerts a power of 250 watts, will he be able to come to class before being marked late? Support your answer.
1
Expert's answer
2017-07-27T13:59:07-0400
The power is
N=Fv
v=N/F=250/500=0.5 m/s.
The maximum distance is
d=vt=0.5(10)=5 m.
It is less than 6m. Thus, the student will not be able to come to class before being marked late.